OpenStudy (anonymous):

Solve. 6(4 – x) + 6x = 24

OpenStudy (gorv):

24-6x+6x=24 =0

OpenStudy (anonymous):

so we started with distrubute 6(4 – x) + 6x = 24 24-6x+6x=24 now combine terms 24-6x=24 -24 -24 -6x=0 -6x/6=0/-6 x=0

OpenStudy (anonymous):

i got 24=24

OpenStudy (anonymous):

nvm o.o

OpenStudy (anonymous):

Yeah it shold be 0

OpenStudy (anonymous):

you can even subsitute it in. 6(4 – 0) + 6(0) = 24 24-0-0=24 24=24